OpenStudy (anonymous):

multiply: 8y^2/7 x 35x/20y

OpenStudy (campbell_st):

\[\frac{8y^2}{7} \times \frac{35x}{20y} = \frac{ 8y^2 \times35x}{7\times20y}\] cancel out the common factors hint.... \[7 \times 20y = 7 \times 5 \times 4\times y\] and \[8y^2 \times 35x = 7\times5\times4\times2\times y \times y \times x\]